OpenStudy (anonymous):

Option 1: A flat salary of $2000 per month. Option 2: A $500 stipend plus 10% of her monthly sales. What would her monthly sales need to be to make Option 2 the better choice? a. Her monthly sales must be at least $1500. b. Her monthly sales must be less than $15000. c. Option 2 will never be a better choice. d. Her monthly sales must be more than $15000.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

let her sale be x we can find out in which time option 2 is higher than 1 \[2000\le 500+0.1x\] so \[x \ge 1500\]

OpenStudy (anonymous):

oh \[x \ge 15000\]

OpenStudy (anonymous):

thanks for the explanation @AndrewNJ :)